OpenStudy (anonymous):

A sandbag was thrown downward from a building . The function f(t)=-16t^2-32+384 shows the height f(t), in feet of the sandbag after t seconds. Part A. Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. Part B. Complete the square of the expression for f(x) to determine the vertex of the graph of f(x) . Would this be a maximum or minimum on the graph? Part C: Use your answer in part B to determine the axis of symmetry for f(x).

OpenStudy (anonymous):

@welshfella do you think you could help?

OpenStudy (anonymous):

@Hero @mathstudent55 @iGreen

OpenStudy (anonymous):

I think I have some of Part A but I'm not sure if I'm right

OpenStudy (welshfella):

i think the function should be -16t^2 -32t + 384

OpenStudy (anonymous):

Oops yeah I did this for part A:

OpenStudy (welshfella):

this factors to -16(t^2 +2t - 24)

OpenStudy (welshfella):

yes you are right

OpenStudy (anonymous):

What does it mean to use the factors to interpret the meaning of the x-intercept?

OpenStudy (welshfella):

the x -intercepts on the graph of this function represent the time when f(t) = 0 - that is when the sandbag reaches the ground

OpenStudy (anonymous):

Now Part B?

OpenStudy (welshfella):

so thats is when t = 4 or t = -6 (we can ignore the -6)

OpenStudy (anonymous):

Wait, what do I write for B?

OpenStudy (welshfella):

as the coefficient of x is negative the graph is a parabola which opens downwards and it will have a maximum value.

OpenStudy (welshfella):

you can complete the square by first halving the coefficient of x and proceed as follows -16(t^2 +2t - 24) = -16 [(t + 1)^2 - 1 - 24] ( half of the 2 in 2t) = -16 [(t + 1)^2 - 25]

OpenStudy (anonymous):

You lost me there now ._.

OpenStudy (welshfella):

you ok about -16(t^2 + 2t - 24) ?

OpenStudy (anonymous):

Yeah, I know how to get that, but everything else got wayyy too complicated

OpenStudy (anonymous):

do you think you could write it out step by step?

OpenStudy (welshfella):

yes ok the vertex form of a trinomial function is (x - a)^2 + b where a is the x coordinate of the axis of symmetry and b is the vertex

OpenStudy (welshfella):